# A Probabilistic Framework for LLM Hallucination Detection via Belief Tree Propagation

### v1

@@ Abstract
This paper focuses on the task of hallucination detection, which aims to determine the truthfulness of LLM-generated statements. To address this problem, a popular class of methods utilize the LLM’s self-consistencies in its beliefs in a set of logically related augmented statements generated by the LLM, which does not require external knowledge databases and can work with both white-box and black-box LLMs. However, in many existing approaches, the augmented statements tend to be very monotone and unstructured, which makes it difficult to integrate meaningful information from the LLM beliefs in these statements. Also, many methods work with the binarized version of the LLM’s belief, instead of the continuous version, which significantly loses information. To overcome these limitations, in this paper, we propose Belief Tree Propagation (BTProp), a probabilistic framework for LLM hallucination detection. BTProp introduces a belief tree of logically related statements by recursively decomposing a parent statement into child statements with three decomposition strategies, and builds a hidden Markov tree model to integrate the LLM’s belief scores in these statements in a principled way. Experiment results show that our method improves baselines by 3%-9% (evaluated by AUROC and AUC-PR) on multiple hallucination detection benchmarks. Code is available at https://github.com/UCSB-NLP-Chang/BTProp.
@@ 1 Introduction
Large language models (LLMs) have shown impressive performance across various domains. Existing LLMs are recognized for their ability to follow human instruction and generate outputs that are both grammatically correct and contextually relevant. However, a notable limitation is their occasional divergence from factual accuracy and real-world knowledge. This issue, commonly referred to as “hallucination”, significantly undermines the reliability and trustworthiness of LLMs.
Various methods have been developed to detect hallucinated statements generated by LLMs. One naive solution is to directly prompt the LLM to output confidence scores for the truthfulness of the statements (referred as “model beliefs" in this paper and prior work [1, 2, 3]). However, due to the LLM’s limited knowledge and poor calibration of the confidence estimation, such beliefs may be mistaken. To resolve this problem, one important line of methods bases the truthfulness decision on the LLM’s internal knowledge as well as logical consistency[1, 2, 3, 4, 5, 6, 7]. More specifically, these approaches typically involve two steps. First, generate an augmented set of statements that are related to the target statements. The augmented statements can be premises or contradictory statements of the target statements [8, 3], or extended passages in the same topic of the target statements [4].Second, ask the LLM to make correctness judgments for every statement in the augmented set, and then integrate these judgments to determine the truthfulness of the target statement based on the logical relationships between the augmented statements and the target statement.
@@
To intuitively understand the idea behind such methods, consider the simplest case where there are two contradictory statements. If the LLM believes both statements are true, it would lead to a logical inconsistency and indicate that one of the decisions is flawed. By having a large number of augmented statements, we could then gather rich information about the LLM’s beliefs and correct the potential decision mistakes by resolving the logical inconsistencies. Compared to other hallucination detection methods, such as retrieval-augmented methods [9, 10, 11] and decisions based on LLM’s internal states [12, 13, 14], such consistency-based methods do not rely on external datasets and can work with both black-box and white-box LLMs, and thus gains wide popularity in the research community.
With the same intuition, we propose Belief Tree Propagation (BTProp), a probabilistic framework that leverages logical consistency across model’s beliefs for LLM hallucination detection. Specifically, given a statement from the LLM’s generation, BTProp first generates a belief tree of logically connected statements by recursively decomposing a parent statement into child statements. As the example shown in Figure 1, the root node of the belief tree is one sentence generated by the LLM about the freezing point of water from the FELM [15] dataset. Each child node is either supportive or contradictory to its parent node. To diversify the statements and the logical relationships, BTProp introduces three types of decomposition paradigms. After that, BTProp builds a directed graphical model (a “hidden Markov tree") [16, 17] in which the LLM’s confidence scores are modeled as observed variables, and the ground-truth truthfulness of the statements as the corresponding hidden variables. Thus the truthfulness decision is turned into a standard principled probabilistic inference problem. By estimating the emission probabilities and transition probabilities properly, we can do inference in this hidden Markov tree model to correct the potential wrong beliefs of LLMs.
Compared to existing hallucination detection methods [4, 7] that directly sample additional responses from the LLM as augmentations, our method organize more diverse statements into a tree structure to better utilize the model’s beliefs. Furthermore, our method can integrate the model beliefs in a principled and probabilistic manner. Experiment results show that our method achieves state-of-the-art performance, improving the hallucination detection performance (AUROC and AUC-PR) by 3% - 9% on FELM-Science [15] and FactCheckGPT [18] datasets.
@@ 2 Related Work
Hallucination has become a prominent topic in the research of LLMs [19, 20, 12, 21, 22]. In existing literature, hallucination typically refers to the phenomenon where LLMs generate nonfactual, nonsensical, or unfaithful outputs. Hallucinations have been observed in various tasks such as text summarization [23, 24, 25, 26], question answering [21], and machine translation [27, 28, 29, 30]. Prior work [31, 32, 32] characterizes hallucination into two main types: factuality hallucination and faithfulness hallucination. Factuality hallucinations refer to outputs that are inconsistent with real-world facts or cannot be verified. In comparison, faithfulness hallucinations shift to the generated content that deviates from the user’s instruction or user-provided contextual information. In this paper, we mainly focuses on factuality hallucinations and aim to better leverage the LLM’s intrinsic capabilities to detect such nonfactual statements within their generations.
@@
Existing approaches to detect factuality hallucination can be divided into two categories. The retrieval-based methods [33, 34, 35, 36, 11, 37, 38, 10] involve comparing the LLM’s output with information retrieved from a reliable knowledge base to assess its factuality. However, these methods depend heavily on external knowledge bases, making them less flexible.To detect hallucinations when an external knowledge base is inaccessible, another category of methods leverages the model’s intrinsic capabilities. Among these intrinsic methods, some approaches leverage the reasoning ability of LLMs to detect and reduce hallucinations [15, 20] via chain-of-thought prompting [39]. Uncertainty-based methods perform token-level or sentence-level uncertainty quantification [40, 18] and use the uncertainty to predict whether the model’s generation is non-factual. When the model’s parameters and hidden states are accessible, probing these hidden states has also been shown to be effective in detecting hallucinations [12, 14, 13]. The sampled-based methods (or consistency-based methods) [4, 6, 7] sample multiple additional responses to the user prompt and then perform consistency checks between the augmented responses and the original model generation. Inconsistencies are then used to detect hallucinations. Our method also aims to detect hallucinations using the model’s intrinsic capabilities and is closely related to the consistency-based methods. Instead of sampling additional responses and performing unstructured consistency checks, our method generates logically-related statements organized in a tree structure. We further propose a probabilistic framework based on the hidden Markov tree model to check consistency and detect hallucinations.
@@
A growing body of research explores improving LLMs’ factuality using the logical consistency across their beliefs [41, 42, 1, 8, 2, 3, 5]. Starting from an initial text (e.g., a statement requiring truthfulness assessment or a question), these methods identify additional texts that are logically connected to the initial text and organize them in tree or graph structures. The inconsistencies of model’s beliefs on the factual correctness of these texts are resolved based on the inferential relations among these texts. One main limitation is that they treat the beliefs (LM-generated truth values or probabilities) as calibrated, which is generally not the case. Motivated by this, we build a probabilistic framework to integrate the model’s beliefs and figure out the most possible and reliable correctness evaluation of the initial text.
3 Methodology
In this section, we will describe the details of the proposed method, which leverages the inconsistency in model beliefs for hallucination detection.
3.1 Problem Formulation
Given a statement $v$ from the LLM-generated text, e.g., ‘A star’s temperature is determined by the amount of mass and energy it has’, and an LLM, our primary goal is to use the LLM to gauge the correctness of the statement. Many previous works use simple prompting techniques to directly elicit the LLM’s belief of the statement, whose output would typically be in the form of confidence scores between $[0,1]$. For example, one approach is to directly ask the LLM whether it believes this statement to be true. By sampling multiple answers from the LLM, we can obtain the confidence score as the percentage of the answers that say ‘yes’.
Denote $S_{v}$ as LLM’s confidence score regarding statement $v$ reflecting the LLM’s original belief of the statement. Then $S_{v}$ should be close to one if the LLM believes the statement is true, and zero otherwise. However, due to LLM’s limited knowledge and capability, $S_{v}$ does not always align with the actual truthfulness of the statement – the scores could be high for some false statements, and low for true ones. In this paper, we will seek to improve the truthfulness judgment with the LLM over its confidence score $S_{v}$.
3.2 Our Method: An Intuitive Overview
@@
The basic idea of our approach is to construct a belief tree, denoted as $\mathcal{T}$, where the root node is the target statement, each child node is a statement logically related to the parent node, and each edge represents the logical connection between the two nodes. We then obtain the confidence scores of all the nodes and use the logical consistency among the scores to correct any potential mistakes in the scores.
@@
Figure 2 shows an example, where the target statement is ‘A star’s temperature is determined by the amount of mass and energy it has”. This statement is true but assume that the LLM produces a low confidence score, $0.1$, for the statement. To correct this error, we can construct a belief tree $\mathcal{T}$ by generating two child nodes from the target node, where the first, which is true, is entailed by the target statement and the second, which is false, contradicts the target statement. Assume that the LLM correctly assigns a high confidence score to the former and a low score to the latter. In this case, we can easily spot that the belief of the target node is logically inconsistent with those of the child nodes, and that increasing the confidence score of the target node would resolve the logical inconsistency, hence correcting the mistake.
@@
Of course, the example above shows the simplest case where the LLM’s beliefs in the child nodes are all correct, whereas, in reality, the child nodes’ beliefs may as well be wrong. To mitigate these additional errors, we can construct a much larger belief tree with greater depth and many different statements. Assuming the errors in the confidence scores are sporadic, we can then expect to correct most of these errors, in both the target and child nodes, by taking into account all the confidence scores and their logical consistency. An example of such a full belief tree is shown in Figure 3.
As implied by the discussion above, our algorithm consists of two components: ❶ Constructing the belief tree, and ❷ inferring the truthfulness of the target statement by integrating the confidence scores across the belief tree. Section 3.3 will discuss the former and Sections 3.4-3.6 the latter.
@@ 3.3 Belief Tree Construction
The success of our algorithm hinges on the diversity of the generated statements. Hence, we instruct the LLM to choose from three types of statement generation strategies.
@@
Strategy 1: Statement Decomposition. For statements containing multiple facts/claims, we can decompose them into individual sub-statements. Figure 1 shows an example where the target statement (Node 0) contains multiple facts (the atmosphere pressure level and the freezing point of water on Mound Everest), which is then decomposed into two statements, one on the atmosphere pressure level and one on the freezing point of water. Statement decomposition can be achieved by prompting the LLM with in-context examples, as listed in Appendix A.1.
@@
Strategy 2: Supportive and Contradictory Premises. In this strategy, the LLM is instructed to generate a set of premises that are each supportive or contradictory to the parent claim. When verifying the truthfulness of Node 2 in Figure 1 about the freezing point of water, we can prompt the model to generate contradictory premises (Node 3 and Node 4) that implies the limited influence of low pressure on the freezing point of water. By leveraging these generated premises, we can indeed correct the model’s wrong belief on Node 2 since the model’s confidence scores on these generated premises are high in this example.
@@
Strategy 3: Statement Correction. In this strategy, the LLM is instructed to generate what it believes to be corrected versions of the parent statement. The Node 5 in Figure 1 shows one example of the statement correction process, where the statement is about the freezing point of water. Then the corrected statements would be almost the same as the parent statement (Node 2), except that the actual temperature is replaced with alternatives that LLM believes may be true. This strategy is implemented via a three-step prompting process. First, the LLM is prompted to generate a question about the key information in the statement. For example, the key information in Node 2 in Figure 1 is the freezing point. Then, the LLM is instructed to answer the generated question, and multiple answers will be sampled. Finally, the LLM is prompted to generate a corrected statement from the original statement if it is wrong according to each sampled answer. Note that there can be multiple corrected statements because the LLM may produce different answers when sampled multiple times. Also, the corrected statements may include the parent statement itself if the LLM believes the parent statement is a likely answer. More details about this process can be found in Appendix A.1.
@@
To select the most appropriate strategies for each parent claim, we introduce the following routing procedure. First, we always start by asking the LLM to attempt the statement decomposition strategy. If the decomposition returns multiple statements, which indicates that the decomposition is meaningful and successful, we would use these decomposed statements as the child statements. If the decomposition returns only a single statement, indicating that the parent statement cannot be further decomposed, we will then prompt the LLM to select between strategies 2 and 3 with a list of rules and examples. The detailed prompts will be listed in Appendix A.1.
@@
For strategies 2 and 3, since each child statement is independently generated, we only need to determine their individual relationship to the parent statement, rather than the joint relationship. To this end, we leverage an off-the-shelf natural language inference (NLI) model to infer the entailment, neutrality, or contradiction between the statements. For each pair of parent statement $u$ and an individual child statement $v$, we derive two NLI relationships, one by setting $u$ as the premise and $v$ as the hypothesis, and the other with $u$ and $v$ switched. The two NLI outputs are then mapped to the aforementioned logical relationships – (entail, entail) mapped to equivalence; (entail, neutral) mapped to entailment; (neutral, entail) mapped to reverse entailment; and any results containing contradict in either direction mapped to contradiction. As mentioned, if the NLI module returns (neutral, neutral), the corresponding child statement will be discarded.
@@
The last step of the belief tree construction is to estimate the model’s belief (i.e., whether the statement is true) on each node. We adopt the simple prompting method discussed in Section 3.1 to obtain LLM’s confidence scores. Specifically, we directly probe the LLM with the prompt ‘True or False? {target statement}’ and use the next token prediction probabilities of the words ‘True’ and ‘False’ to compute the model confidence. We normalize the prediction logits of the two words to get the confidence score of that statement. The only exception is in the case of statement correction, where the LLM is already instructed to output alternative statements it believes to be true. As a result, we simply set the confidence score of each generated statement to 1. Our empirical analysis shows that this would greatly reduce the number of LLM queries without deteriorating the performance.

### v2

@@ Abstract
We describe Belief Tree Propagation (BTProp), a probabilistic framework for LLM hallucination detection. To judge the truth of a statement, BTProp generates a belief tree by recursively expanding the initial statement into a set of logically related claims, then reasoning globally about the relationships between these claims. BTProp works by constructing a probabilistic model of the LM itself: it reasons jointly about logical relationships between claims and relationships between claim probabilities and LM factuality judgments via probabilistic inference in a “hidden Markov tree”. This method improves over state-of-the-art baselines by 3%-9% (evaluated by AUROC and AUC-PR) on multiple hallucination detection benchmarks. Code is available at https://github.com/UCSB-NLP-Chang/BTProp.
A Probabilistic Framework for LLM Hallucination Detection
via Belief Tree Propagation
Bairu Hou1thanks: Correspondence to: Bairu Hou $<$bairu@ucsb.edu$>$. Yang Zhang2 Jacob Andreas3 Shiyu Chang1 1UC Santa Barbara  2MIT-IBM Watson AI Lab  3MIT CSAIL
@@ 1 Introduction
Current large language models (LLMs) often produce factually incorrect statements (sometimes referred to as “hallucinations”). One popular approach to detecting hallucinations is to prompt LLMs to assign probabilities to the correctness of their own outputs (sometimes referred to as model “beliefs"; Mitchell et al., 2022; Hase et al., 2023; Kassner et al., 2023). However, these beliefs are themselves frequently incorrect, miscalibrated, or inconsistent with each other. To improve the accuracy of LM factuality judgments, a large body of work adjusts the assessed probability of a target claim by reasoning about sets of related statements Manakul et al. (2023); Akyürek et al. (2024); Cao et al. (2023); Mündler et al. (2023). These approaches typically involve two steps: first, the LLM is prompted to generate an augmented set of statements that support or contradict the target statement, possibly with probabilities attached Jung et al. (2022); second, these claims are combined via logical or probabilistic inference. As a simple example, if an LM assigns high probability to a target statement, but even higher probability to a contradictory statement, it may be reasonably inferred that the target statement is actually incorrect.
@@
One key limitation of current methods is that they do not model noise or uncertainty in the LM generation process itself. LLMs’ “beliefs” are often poorly calibrated – even statements that models judge to be true with 100% probability are sometimes false. How can we produce reliable judgments about sets of related statements when even their assessments of individual statements are unreliable?
To do so, we propose Belief Tree Propagation (BTProp), a probabilistic method for improving the accuracy LLM “beliefs” with applications to hallucination detection. Specifically, given an initial statement generated by an LLM, BTProp first generates a belief tree of recursively generated claims that are logically related ot the initial statement. An example is shown in Figure 1, in which the root node of the belief tree is one sentence generated by the LLM about the freezing point of water, and each child node supports or contradicts its parent. Next, BTProp interprets this belief tree as a directed graphical model (a “hidden Markov tree"; Crouse et al., 1998; Durand et al., 2004) in which the LLM’s confidence scores are modeled as observed variables, and the ground-truth truthfulness of the statements as the corresponding hidden variables. Thus, the problem of assigning probabilities to statements can be reduced to probabilistic inference in a well-studied model family. Here, inference makes it possible to integrate knowledge about the (logical) structure of related statements, and the (probabilistic) relationship between statements and noisy LM outputs. Experiment results show that our method achieves state-of-the-art performance, improving the hallucination detection performance (AUROC and AUC-PR) by 3%–9% on FELM-Science Chen et al. (2023b) and FactCheckGPT Fadeeva et al. (2024) datasets relative to state-of-the-art baselines.
@@ 2 Related Work
Hallucination has become a prominent topic in the research of LLMs Liu et al. (2022); Dhuliawala et al. (2023); Azaria and Mitchell (2023); Li et al. (2023); Zhang et al. (2023a). Prior work Ji et al. (2023); Huang et al. (2023) characterizes hallucination into two main types: factuality hallucination and faithfulness hallucination. Factuality hallucinations refer to outputs that are inconsistent with real-world facts. In comparison, faithfulness hallucinations shift to the generated content that deviates from the user’s instruction or user-provided contextual information. In this paper, we mainly focuses on factuality hallucinations and aim to better leverage the LLM’s intrinsic capabilities to detect such nonfactual statements within their generations.
@@
To detect hallucinations in LLM generation, the retrieval-based methods Shuster et al. (2021); Min et al. (2023); Chern et al. (2023); Semnani et al. (2023); Huo et al. (2023); Zhang et al. (2024) compare the LLM’s output with information from a reliable knowledge base. When an external knowledge base is inaccessible, another category of methods rely on the model’s intrinsic capabilities. Among them, some approaches leverage the reasoning ability of LLMs to detect and reduce hallucinations Chen et al. (2023b); Dhuliawala et al. (2023) via chain-of-thought prompting Wei et al. (2022). Uncertainty-based methods perform token-level or sentence-level uncertainty quantification Varshney et al. (2023); Fadeeva et al. (2024) to predict the factuality of model’s generation. Probing the model’s hidden states can also help detect hallucinations Azaria and Mitchell (2023); Chen et al. (2023a); Su et al. (2024). The sampled-based methods (or consistency-based methods) Manakul et al. (2023); Cao et al. (2023); Mündler et al. (2023) sample multiple responses and then perform consistency checks between these responses and the original model generation. Inconsistencies are then used to detect hallucinations. Our method is closely related to the consistency-based methods. Instead of sampling additional responses and performing unstructured consistency checks, our method generates logically-related statements organized in a tree structure. We further propose a probabilistic framework based on the hidden Markov tree model to check consistency and detect hallucinations.
@@
A growing body of research explores improving LLMs’ factuality using the logical consistency across their beliefs Dalvi et al. (2021); Tafjord et al. (2022); Mitchell et al. (2022); Jung et al. (2022); Hase et al. (2023); Kassner et al. (2023); Akyürek et al. (2024). Starting from an initial text (e.g., a statement requiring truthfulness assessment or a question), these methods identify additional texts that are logically connected to the initial text and organize them in tree or graph structures. The inconsistencies of model’s beliefs on the factual correctness of these texts are resolved based on the inferential relations among these texts. One main limitation is that they treat the beliefs (LM-generated truth values or probabilities) as calibrated, which is generally not the case. Motivated by this, we build a probabilistic framework to integrate the model’s beliefs and figure out the most possible and reliable correctness evaluation of the initial text.
3 Method
Given a statement $v$ from the LLM-generated text, e.g., A star’s temperature is determined by the amount of mass and energy it has, and an LLM, we aim to use the LLM to gauge the correctness of the statement. Many previous approaches use simple prompting techniques to directly elicit the LLM’s “belief”, or assessed probability of the statement, as a confidence score in $[0,1]$. For example, one approach is to directly query the LLM for whether this is true. By sampling multiple answers from the LLM, we can obtain the confidence score as the fraction of answers that say yes (or some paraphrase thereof). Crucially, these scores may be wrong in several ways: in addition to being factually incorrect, they may be miscalibrated (reflecting an inappropriate degree of certainty or uncertainty) or inconsistent (e.g., assigning high probability to mutually incompatible statements, or different probabilities to logically equivalent statements). Below, we describe a method that addresses both categories of error by reasoning jointly about relationships between the truth values of related statements and LLM confidence scores.
3.1 BTProp: An Intuitive Overview
@@
The basic idea of our approach is to construct a belief tree, denoted $\mathcal{T}$, in which the root node is the target statement, each child node is a statement logically related to the parent node, and each edge represents the logical relationship between two nodes. We then obtain the confidence scores of all the nodes and use the logical consistency among them to correct any potential mistakes in the scores.
@@
Figure 2 shows an example, where the target statement is A star’s temperature is determined by the amount of mass and energy it has. This statement is true but assume that the LLM produces a low confidence score, $0.1$, for the statement. To correct this error, we can construct a belief tree $\mathcal{T}$ by generating two child nodes from the target node, where the first, which is true, is entailed by the target statement and the second, which is false, contradicts the target statement. Assume that the LLM correctly assigns a high confidence score to the former and a low score to the latter. In this case, we can easily recognize that the belief of the target node is logically inconsistent with those of the child nodes, and that increasing the confidence score of the target node would resolve the logical inconsistency, thereby correcting the mistake.
@@
The example above shows a simple case in which LLM beliefs about child nodes are all correct. In reality, these child nodes’ beliefs may themselves be incorrect. To mitigate these additional errors, we can construct a much larger belief tree with greater depth and many different statements. Assuming the errors in the confidence scores are sporadic, we can then expect to correct most of these errors, in both the target and child nodes, by taking into account all the confidence scores and their logical consistency. An example of such a full belief tree is shown in Figure 3.
As implied by the discussion above, our algorithm consists of two components: ❶ Constructing the belief tree, and ❷ inferring the truthfulness of the target statement by integrating the confidence scores across the belief tree. Section 3.2 will discuss the former and Sections 3.3-3.5 the latter.
@@ 3.2 Belief Tree Construction
As will be described below, BTProp is general, and compatible with any method for generating belief tree nodes from parent statements. We explore three specific strategies for statement generation:
@@
Strategy 1: Statement Decomposition. For statements containing multiple facts/claims, we decompose them into individual sub-statements. Figure 1 shows an example where the target statement (Node 0) contains multiple facts (the atmosphere pressure level and the freezing point of water on Mound Everest), It is then decomposed into two statements, one on the atmosphere pressure level and one on the freezing point of water. Statement decomposition can be achieved by prompting the LLM with in-context examples, as listed in Appendix A.7.
@@
Strategy 2: Supportive and Contradictory Premises. We prompt the LLM to generate a set of premises that are supportive or contradictory to the parent claim. When verifying the truthfulness of Node 2 in Figure 1 about the freezing point of water, we can prompt the model to generate contradictory premises (Node 3 and Node 4) that implies the limited influence of low pressure on the freezing point of water. By leveraging these generated premises, we can indeed correct the model’s wrong belief on Node 2 since the model’s confidence scores on these generated premises are high in this example.
@@
Strategy 3: Statement Correction. In this strategy, the LLM is instructed to generate what it believes to be corrected versions of the parent statement. The Node 5 in Figure 1 shows one example of the statement correction process, where the statement is about the freezing point of water. Then the corrected statements would be almost the same as the parent statement (Node 2), except that the actual temperature is replaced with alternatives that LLM believes may be true. This strategy is implemented via a three-step prompting process. First, the LLM is prompted to generate a question about the key information in the statement (e.g., the freezing point in Node 2 in Figure 1). Then, the we sample answers from the LLM. Finally, the LLM is prompted to generate a corrected statement from the original statement if it is wrong according to each sampled answer. Note that there can be multiple corrected statements because the LLM may produce different answers when sampled multiple times. Also, the corrected statements may include the parent statement itself if the LLM believes the parent statement is a likely answer. More details about this process can be found in Appendix A.7.
@@
Our method select the most appropriate strategies for each parent claim as follows. First, LLM first attempt the statement decomposition strategy. If the decomposition returns multiple statements, which indicates that the decomposition is meaningful, we would use them as the child statements. If only a single statement is returned, indicating that the parent statement cannot be further decomposed, we will prompt the LLM to select between strategies 2 and 3 with a list of rules and examples. The detailed prompts are in Appendix A.1.
@@
For strategies 2 and 3, since each child statement is independently generated, we only need to determine their individual relationship to the parent statement, rather than the joint relationship. To this end, we leverage an off-the-shelf natural language inference (NLI) model to infer the entailment, neutrality, or contradiction between the statements. For each pair of parent statement $u$ and an individual child statement $v$, we derive two NLI relationships, one by setting $u$ as the premise and $v$ as the hypothesis, and the other with $u$ and $v$ switched. The two NLI outputs are then mapped to the aforementioned logical relationships – (entail, entail) mapped to equivalence; (entail, neutral) mapped to entailment; (neutral, entail) mapped to reverse entailment; and any results containing contradict in either direction mapped to contradiction. If the NLI module returns (neutral, neutral), the corresponding child statement will be discarded.
@@
The last step of the belief tree construction is to estimate the model’s belief (i.e., whether the statement is true) on each node. Specifically, we directly probe the LLM with the prompt ‘True or False? {target statement}’ and use the next token prediction probabilities of the words ‘True’ and ‘False’ to compute the model confidence. We normalize the prediction logits of the two words to get the confidence score of that statement. The only exception is in the case of statement correction, where the LLM is already instructed to output alternative statements it believes to be true. As a result, we simply set the confidence score of each generated statement to 1. Our empirical analysis shows that this would greatly reduce the number of LLM queries without deteriorating the performance.